# Deep Time Series Models: A Comprehensive Survey and Benchmark

### v1

@@ 4.4 Graph Neural Networks
Analyzing multivariate time series data is often challenging due to the complex and often non-linear correlations between variables. To address this challenge, Graph neural networks (GNNs) [136, 137] have been widely adopted in time series analysis. By modeling multivariate data as a spatiotemporal graph, where each node represents a variable, GNNs can extract relationships among neighboring nodes and capture the temporal evolution of node attributes over time, thereby providing a robust framework for understanding the underlying dynamics of multivariate time series.
The core goal of GNN architecture is to model the underlying topological relations in multivariate data, therefore existing GNN-based works can be roughly divided into two categories based on whether graph structure is part of the input into the model. DCRNN [138] models the spatial dependency of traffic as a diffusion process on a directed graph and uses diffusion convolution to capture the spatial dependency, alongside a recurrent neural network to capture temporal dynamics. Similarly, STGCN [139] integrates graph convolutional networks to model the spatial dependencies among traffic sensors with temporal convolutions to capture the temporal dependencies in the traffic time series data. Graph WaveNet [140] combines graph convolution with dilated casual convolution and learns an adaptive dependency matrix through node embedding, enabling the model to automatically capture hidden spatial dependencies in spatial-temporal graph data. Similarly, AGCRN [141] enhances the traditional graph convolutional network with node adaptive parameter learning and data-adaptive graph generation modules, allowing for the automatic capture of spatial and temporal correlations without a pre-defined graph structure. MTGNN [142] introduces a graph learning layer to adaptively learn the graph adjacency matrix, thereby capturing hidden relationships among multivariate time series data. STFGNN [143] employs a Spatial-Temporal Fusion Graph Neural Network with a generated temporal graph to learn localized spatial-temporal heterogeneity and global spatial-temporal homogeneity. StemGNN [28] leverages the advantages of both the Graph Fourier Transform (GFT) and the Discrete Fourier Transform (DFT), modeling multivariate time series in the spectral domain.
@@ 4.5 Transformers
In the view of the great success in the field of natural language processing [16, 144, 145, 146, 147] and computer vision [19, 148, 149, 150], Transformers have also emerged as a powerful backbone for time series analysis. Benefiting from the self-attention mechanism [151], Transformer-based models can capture long-term temporal dependencies and complex multivariate correlations. As overviewed in Figure 8, existing Transformer-based time series models can be categorized based on the granularity of representation used in the attention mechanism, namely point-wise, patch-wise, and series-wise approaches.
@@ 4.5.1 Point-wise Dependency
Due to the serial nature of time series, most existing Transformer-based works use a point-wise representation of time series data and apply attention mechanisms to capture the correlations among different time points. Among these point-wise modeling approaches, Data Embedding is a crucial component that maps the value of time series data to a high-dimensional representation. Given time series with corresponding time stamp information , where is the variate number and is the types of time stamps, the embedding module can be summarized as follow:
@@
To better apply the Transformer architecture to the time series domain, researchers have explored two aspects: designing pre-processing modules and modifying the attention mechanisms. As we have discussed in Section 3.1, past RevIN [40] and Stationary [41] achieved superior performance by introducing Normalization and De-Normalization modules before and after the vanilla Transformer. Besides, Stationary [41] further proposes De-stationary Attention to avoid the over-stationarization problem.
Given that the canonical attention approach leads to a quadratic computational complexity, numerous efficient Transformers [21, 22, 153, 86] have been proposed to mitigate the complexity caused by point-wise modeling, which is summarized in Table II. LogSparse [152] proposes Convolutional Self-Attention to replace canonical attention by employing causal convolutions to produce queries and keys in the self-attention layer. Informer [21] introduces a Query Sparsity Measurement, where a larger value indicates a higher chance to contain the dominant information in self-attention. Based on the proposed sparsity measurement, it further designs a ProbSparse self-attention only using top queries with the biggest measurement results, which can reduce the quadratic complexity in time and memory. Pyraformer [153] constructs a multi-resolution -ary tree and develops a Pyramidal Attention Mechanism, in which every node can only attend to its neighboring, adjacent, and children nodes. With the calculated mask for attention, Pyraformer can capture both short- and long-temporal dependencies with linear time and space complexity.
@@ 4.5.2 Patch-wise Dependency
Patch-based architectures play a crucial role in the Transformer models for both Natural Language Processing (NLP) [16] and Computer Vision (CV) [19]. Since point-wise representations are insufficient to capture local semantic information in temporal data, several studies [23, 154, 155] have been devoted to exploring patch-level temporal dependencies within time series data.
Pioneer work Autoformer [22] proposes an Auto-Correlation Mechanism, which captures the series-wise dependencies of time series to replace canonical point-wise self-attention. Based on the stochastic process theory [156], Auto-Correlation utilizes the Fast Fourier Transform to discover the time-delay similarities between different sub-series. A time delay module is further proposed to aggregate the similar sub-series from underlying periods instead of the relation between scattered points, which firstly explores the sub-series level modeling in Transformer-based models.
Different from modifying the attention mechanism, most recent works utilize patch-wise representations of time series data and perform a self-attention mechanism to capture patch-wise dependencies [23, 54, 157]. PatchTST [23] and follow-up works split time series into a sequence of overlapped patches and embed each patch following:
Assume , is patch length and the corresponding number of patches split, and denotes the -th patch with sequence length . The patches are mapped to the latent space through a temporal linear projection and a learnable position embedding .
Based on the vanilla attention mechanism, PatchTST [23] learns the patch-wise dependencies. Going beyond PatchTST, recent Pathformer [158] proposes a multi-scale Transformer-based model with adaptive pathways. Based on the patch division of different scales, the adaptive pathways select the patch sizes with the top weights generated by the router to capture multi-scale characteristics.
The success of PatchTST also benefits from channel-independence design, where each temporal patch-level token only contains information from a single series. In addition to capturing the patch-level temporal dependencies within one single series, recent approaches [54, 157] have endeavored to capture interdependencies among patches from different variables over time. Crossformer [29] introduces a Two-Stage Attention layer containing a Cross-Time Stage and a Cross-Dimension Stage to efficiently capture the cross-time and cross-variate dependencies between each patch token. For the obtained embedded vector , the overall attention stage can be described as follow:
@@ 4.5.3 Series-wise Dependency
Further expanding the receptive field, there are also some works that attempt to use the tokenization of the whole time series to capture inter-series dependencies. iTransformer [30] introduce to multivariate data, and for -th variable , it can be simply formulated as follows:
@@ 5 Time Series Library
Time series analysis has emerged as an important research area, attracting significant attention from both academia and industry. Recently, extensive exploration of deep learning based methods for time series analysis has resulted in significant advances. However, the issue of fair benchmarking poses a pressing challenge in this domain. The absence of fair, rational, and comprehensive benchmarks can lead to biased comparisons between different methods and hinder accurate evaluation of their effectiveness, potentially inflating domain advances or hindering practical applications. This presents a substantial obstacle to understanding advances and fostering robust development within the field.
In the domain of time series analysis, several benchmarks have been proposed, such as DGCRN [160], LibCity [161], DL-Traff [162], TS-bench [163], and BasicTS [14]. More specifically, Autoformer [22] proposed a standard long-term forecasting benchmark covering different practical applications. Further, to verify the generality of different time series analysis models, TimesNet [3] builds a more comprehensive model generalization benchmark covering five mainstream time series analysis tasks. However, these benchmarks typically have some limitations. One issue with current time series benchmarks is their limited coverage of time series analysis tasks and specific domains, which limits their practical applications. Moreover, these benchmarks often fail to provide detailed discussions and comprehensive summaries of task types, model architectures, and specific baseline methods. As a result, they do not effectively guide the design of more efficient time series analysis methods or drive further development in the field.
To effectively address these issues, we introduce and implement Time Series Library (TSLib), a benchmark for fair and comprehensive comparing and evaluating the performance of deep time series models across various time series analysis tasks. As shown in Figure 9, TSLib encompasses a unified model experiment pipeline, standardized evaluation protocols, extensive and diverse real-world datasets, mainstream and advanced time series analysis models, and unified experimental validation and analysis process.
In our Time Series Library, we meticulously followed the official codes and implemented 24 widely used and advanced deep time series analysis models. These models are derived from four canonical deep learning architectures. Users can choose from these models based on their specific practical usage scenarios. The code is available at https://github.com/thuml/Time-Series-Library.
As follows, we will provide a detailed description of our TSLib, including the design and implementation principles (Section 5.1), the evaluation protocols and metrics (Section 5.2), the dataset descriptions (Section 5.3), and the main results of models with different architectures (Section 5.4).
@@ 5.1 Design and Implementation Principle
TSLib is designed based on the well-established factory pattern and implements a unified interface between data and model objects, thus enabling a clear separation between deep model creation and usage, promoting modularity and flexibility. By loading different data and model objects and combining specific task heads during model training, TSLib enables different datasets and models to be shared and extended, allowing easy switching between various time series analysis tasks. These design and implementation principles provide enhanced flexibility and scalability for our TSLib. Furthermore, as illustrated in Figure 9, TSLib introduces a unified experimental pipeline covering the overall process of the model training and evaluation, which includes data source, data processing, model training and analysis, and model performance evaluation.
@@
Data Source Our TSLib provides extensive support for a wide range of diverse and multi-type datasets in a variety of formats, including ”.csv”, ”.npz”, ”.txt”, etc. As shown in Figure 9 and Table III, TSLib currently supports more than 30 datasets with different sampled frequencies across four mainstream time series analysis tasks, all derived from real-world scenarios in domains such as energy, transportation, economics, weather, and medicine, etc. Moreover, TSLib excels in scalability, allowing for the effortless integration of new data sources of different data types.
Data Processing Data processing plays a pivotal role in guaranteeing stable training within the realm of time series analysis. Within the Time Series Library, a multitude of data processing steps are conducted, including time window splitting, data batch generation, etc. Subsequently, the raw data is partitioned into separate sets for training, validation, and testing purposes, enabling streamlined model training and equitable comparisons. These steps serve as indispensable prerequisites for attaining precise and dependable results across a range of diverse time series analysis tasks.
@@
Moreover, our TSLib provides additional support for numerous crucial and effective data processing strategies based on different model design principles [39], [40], [41], [45] to enhance model performance and training efficiency. We encapsulate these various design strategies within our basic data processing layer, encompassing techniques such as data normalization, time-frequency decomposition, Fourier analysis, and more. When utilizing TSLib, users have the flexibility to select these strategies to improve training effect based on their specific requirements and objectives.
Model Training and Analysis After the data processing phase, the raw time series data is transformed into the desired format for model training. Model training forms the crux of the entire experiment pipeline, where we fine-tune the model parameters based on the input to predict the output with minimal error. Our primary goal during model training is to obtain the best possible trainable parameters that result in a significant improvement in model performance. Each model has its own unique design and training objective. The model analysis procedure is to determine the optimal model parameters by comparing the correlation between the training and validation losses. Our TSLib includes complete log printing and result storage functions record and evaluate the training process. By employing rational model analysis techniques, we can efficiently obtain models with superior performance and stronger generalization.
Performance Evaluation Model evaluation is a crucial step in verifying the effectiveness and generalization of trained time series models. It involves model prediction and performance evaluation, providing insights into the efficacy of the trained model. In TSLib, we provide evaluation support for four mainstream time series analysis tasks: classification, imputation, forecasting (long-term or short-term), and anomaly detection. Each task comes with its specific evaluation metric, enabling a comprehensive assessment of the performance of models. These metrics play a crucial role in determining the effectiveness of the trained model and its suitability for the intended task.
| Task | Dataset | Dimension | Length | Domain | Size |
| Classification | EthanolConcentration | 3 | 1,751 | Alcohol Industry | 20.3 MB |
| FaceDetection | 144 | 62 | Face (250 Hz) | 789.1 MB | |
| Handwriting | 3 | 152 | Motion | 3.9 MB | |
| Heartbeat | 61 | 405 | Health (0.061 secs) | 87.8 MB | |
| JapaneseVowels | 12 | 29 | Voice | 1.1 MB | |
| PEMS-SF | 963 | 144 | Transportation (1 day) | 420.1 MB | |
| SelfRegulationSCP1 | 6 | 896 | Health (256 Hz) | 17.8 MB | |
| SelfRegulationSCP2 | 7 | 1,152 | Health (256 Hz) | 17.7 MB | |
| SpokenArabicDigits | 13 | 93 | Voice (11025 Hz) | 37.6 MB | |
| UWaveGestureLibrary | 3 | 315 | Gesture | 3.4 MB | |
| Imputation | ETTh1, ETTh2 | 7 | 17,420 | Electricity (1 hour) | 10.4 MB |
| ETTm1, ETTm2 | 7 | 69,680 | Electricity (15 mins) | 2.6 MB | |
| Electricity | 321 | 26,304 | Electricity (1 hour) | 95.6 MB | |
| Weather | 21 | 52,696 | Environment (10 mins) | 7.2 MB | |
| Long-term Forecasting | ETTh1, ETTh2 | 7 | 17,420 | Electricity (1 hour) | 10.4 MB |
| ETTm1, ETTm2 | 7 | 69,680 | Electricity (15 mins) | 2.6 MB | |
| Electricity | 321 | 26,304 | Electricity (1 hour) | 95.6 MB | |
| Weather | 21 | 52,696 | Environment (10 mins) | 7.2 MB | |
| Traffic | 862 | 17,544 | Transportation (1 hour) | 136.5 MB | |
| Exchange | 8 | 7,588 | Economic (1 day) | 623 KB | |
| ILI | 7 | 966 | Health (1 week) | 66 KB | |
| Short-term Forecasting | M4-Yearly | 1 | 6 | Demographic | 589.5 MB |
| M4-Quarterly | 1 | 8 | Finance | ||
| M4-Monthly | 1 | 18 | Industry | ||
| M4-Weakly | 1 | 13 | Macro | ||
| M4-Daily | 1 | 14 | Micro | ||
| M4-Hourly | 1 | 48 | Other | ||
| Anomaly Detection | SMD | 38 | 100 | Industry (1 min) | 436.4 MB |
| MSL | 55 | 100 | Industry (1 min) | 58.2 MB | |
| SMAP | 25 | 100 | Industry (1 min) | 113.0 MB | |
| SwaT | 51 | 100 | Industry (1 min) | 903.2 MB | |
| PSM | 25 | 100 | Industry (1 min) | 107.1 MB |
5.2 Evaluation Protocols
In order to conduct a fair and comprehensive model performance verification, our Time Series Library is designed to provide standardized evaluation protocols for four mainstream time series analysis tasks following [3]. The primary goal of these standardized and unified evaluation protocols is to quantify the effectiveness of different time series analysis methods with varying architectures. Additionally, they provide valuable insights into the strengths and limitations of different methods across diverse time series analysis tasks. By establishing these standardized evaluation protocols, we aim to promote fair comparisons between different methods and improve our understanding of their performance in various time series analysis scenarios.

### v2

@@ 4.4 Graph Neural Networks
Analyzing multivariate time series data is often challenging due to the complicated and often nonlinear correlations between variables. To address this, Graph neural networks (GNNs) [137, 138] have been widely adopted in time series analysis. As depicted in Figure 6, by modeling multivariate data as a spatiotemporal graph, where each node represents a variable, GNNs can extract relationships among neighboring nodes and capture the temporal evolution of node attributes over time, providing a framework for understanding the underlying dynamics of multivariate time series.
The primary goal of GNN architecture is to model the latent topological relations in multivariate data. Accordingly, existing GNN-based works can be broadly categorized into two types: those that rely on a pre-defined graph structure as input and those that learn the graph structure dynamically. DCRNN [139] models the spatial dependency of traffic as a diffusion process on a directed graph and uses diffusion convolution to capture the spatial dependency, alongside a recurrent neural network to capture temporal dynamics. Similarly, STGCN [140] integrates graph convolutional networks to model the spatial dependencies among traffic sensors with temporal convolutions to capture the temporal dependencies in the traffic time series data. Graph WaveNet [141] combines graph convolution with dilated causal convolution and learns an adaptive dependency matrix through node embedding, enabling the model to automatically capture hidden spatial dependencies in spatial-temporal graph data. Similarly, AGCRN [142] enhances the traditional graph convolutional network with node adaptive parameter learning and data-adaptive graph generation modules, allowing for the automatic capture of spatial and temporal correlations without a pre-defined graph structure. MTGNN [143] introduces a graph learning layer to adaptively learn the graph adjacency matrix, thereby capturing hidden relationships among multivariate time series data. STFGNN [144] employs a Spatial-Temporal Fusion Graph Neural Network with a generated temporal graph to learn localized spatial-temporal heterogeneity and global spatial-temporal homogeneity. StemGNN [87] leverages the advantages of both the Graph Fourier Transform (GFT) and the Discrete Fourier Transform (DFT), modeling multivariate time series in the spectral domain.
@@ 4.5 Transformers
In view of the great success in the field of natural language processing [16, 145, 146, 147, 148] and computer vision [19, 149, 150, 151], Transformers have emerged as a powerful backbone for time series analysis. Benefiting from the self-attention mechanism [152], Transformer-based models excel at capturing long-term temporal dependencies and entangled multivariate correlations. As overviewed in Figure 7, existing Transformer-based time series models can be categorized based on the granularity of representation used in the attention mechanism, namely point-wise, patch-wise, and series-wise approaches.
@@ 4.5.1 Point-wise Dependency
Due to the serial nature of time series, most existing Transformer-based works learn point-wise representations and apply attention mechanisms to capture the correlations among different time points. Among these point-wise modeling approaches, Data Embedding stands as a crucial component, mapping raw time series values into a high-dimensional representation. Given time series with corresponding time stamp information , where is the variate number and is the types of time stamps, the embedding module can be summarized as:
@@
To optimize the application of Transformer architecture to the time series domain, existing literature primarily falls into two categories: designing pre-processing modules and modifying the attention mechanisms. As previously discussed in Section 3.1, techniques like RevIN [37] and Stationary [38] have achieved superior performance by integrating Normalization and De-Normalization modules before and after the Transformer blocks. Notably, Stationary [38] further introduces De-stationary Attention to mitigate the problem of over-stationarization.
Given that canonical attention leads to a quadratic computational complexity, various efficient attentions [21, 22, 154, 83] have been proposed to mitigate the complexity arising from point-wise modeling, which is summarized in Table II. LogSparse [153] proposes Convolutional Self-Attention to replace canonical attention by employing causal convolutions to produce queries and keys in the self-attention layer. Informer [21] introduces a Query Sparsity Measurement, where a larger value indicates a higher chance of containing the dominant information in self-attention. Building on this, it designs ProbSparse self-attention, which selectively uses only the top queries with the highest measurement results, thereby significantly reducing quadratic complexity in both time and memory. Pyraformer [154] constructs a multi-resolution tree and develops a Pyramidal Attention Mechanism, in which every node can only attend to its neighboring, adjacent, and children nodes. With a calculated mask for attention, Pyraformer captures both short- and long-temporal dependencies with linear computational complexity.
@@ 4.5.2 Patch-wise Dependency
Patch-based architectures play a crucial role in the Transformer models for both Natural Language Processing (NLP) [16] and Computer Vision (CV) [19]. Since point-wise representations often fall short in capturing local semantic information in temporal data, several studies [23, 156, 157] have been devoted to exploring patch-level temporal dependencies in time series data.
Pioneer work Autoformer [22] proposes an Auto-Correlation Mechanism to replace canonical point-wise self-attention, enabling the model to learn series-wise dependencies. Based on the stochastic process theory [158], Auto-Correlation utilizes the Fast Fourier Transform to discover the time-delay similarities between different sub-series. A time delay module is further proposed to aggregate similar sub-series from underlying periods instead of the relation between scattered points, which firstly explores the sub-series level modeling in Transformer-based models.
Different from modifying the attention mechanism, most recent works utilize patch-wise representations of the data and perform a canonical self-attention mechanism to capture patch-wise dependencies [23, 52, 159]. PatchTST [23] and its subsequent works split time series into a sequence of overlapped patches and embed each patch following:
Let denote the patch length and the corresponding number of patches, and is the -th patch of length . The patches are mapped to the latent space through a temporal linear projection and a learnable position embedding . Based on the vanilla attention mechanism, PatchTST [23] learns the patch-wise dependencies. Going beyond PatchTST, Pathformer [160] introduces a multi-scale Transformer that employs adaptive pathways to dynamically select patch sizes, thereby learning multi-scale temporal representations.
The success of PatchTST also benefits from its channel-independence design, where each temporal patch-level token encapsulates information from only a single series. In addition to capturing the patch-level temporal dependencies within one single series, recent approaches [52, 159] have endeavored to capture interdependencies among patches from different variables over time. Crossformer [155] introduces a Two-Stage Attention layer containing a Cross-Time Stage and a Cross-Dimension Stage to efficiently capture the cross-time and cross-variate dependencies between each patch token. For the obtained embedded vector , the overall attention stage can be described as follow:
@@ 4.5.3 Series-wise Dependency
Further expanding the receptive field, recent works tokenize the entire time series to capture inter-series dependencies. iTransformer [39] introduces for multivariate data and for -th variable , it can be formulated as:
@@ 5 Time Series Library
Recently, extensive exploration of deep time series methods has resulted in significant advances. Given such booming development, the issue of fair benchmarking poses a pressing challenge for future research. Thus, focusing on time series analysis, several benchmarks have been proposed, such as DGCRN [162], LibCity [163], DL-Traff [164], TS-bench [165], and BasicTS [14]. Besides, Autoformer [22] proposed a standard long-term forecasting benchmark covering different practical applications. Further, to verify the generality of different time series analysis models, TimesNet [3] builds a more comprehensive model generalization benchmark covering five mainstream time series analysis tasks. However, these benchmarks typically have some limitations. One issue with current time series benchmarks is their limited coverage of time series analysis tasks and specific domains, which limits their practical applications. Moreover, these benchmarks often fail to provide detailed discussions and comprehensive summaries of task types, model architectures, and specific baseline methods. As a result, they do not effectively guide the design of more efficient time series methods or drive further development in the field.
To address these challenges effectively, we introduce Time Series Library (TSLib), a benchmark designed for the fair and comprehensive evaluation of deep time series models across a wide range of analysis tasks. As shown in Figure 8, TSLib provides a unified experimental framework that includes standardized evaluation protocols, a rich and diverse collection of real-world datasets, state-of-the-art and widely-used time series models, and a streamlined process for experimental validation and analysis. Although numerous time series libraries have been proposed in recent years, TSLib stands out by including advanced models and supporting a wide range of time-series analysis tasks, which have earned widespread recognition from the community as shown in Table III. Specifically, we meticulously followed the official codes and implemented 30 widely used and advanced deep time series models. Users can choose from these models based on their specific practical usage scenarios. The code is available at https://github.com/thuml/Time-Series-Library.
Library |
Models |
Task |
Stars |
tslearn∗ [166] |
10 |
Clustering, Classification, Regression |
3k |
StatsForecast∗ [167] |
31 |
Forecasting, Anomaly Detection |
4.4k |
NeuralForecast [168] |
33 |
Forecasting |
3.6k |
GluonTS [169] |
25 |
Forecasting, Anomaly Detection |
4.9k |
| TSLib (This Survey) | 30 |
Forecasting, Imputation |
9.4k |
Classification, Anomaly Detection |
This section provides an overview of TSLib. We will introduce design and implementation principles of TSLib in Section 5.1, followed by a brief description of default evaluation metrics in Section 5.2, supported datasets in Section 5.3, and experimental settings in Section 5.4. Finally, comparisons of various models are included in Section 5.5.
@@ 5.1 Design and Implementation Principle
TSLib is designed based on the well-established factory pattern and implements a unified interface between data, model, and experiment objects, thus enabling a clear separation between deep model creation and usage, promoting modularity and flexibility. By loading different data and model objects and combining task-specific heads, TSLib enables different datasets and models to be shared and extended, allowing easy switching between various datasets, models, and analysis tasks. Here are the key features.
@@
Diverse Data Formats TSLib natively supports a wide range of datasets in a variety of formats, including “.csv”, “.npz”, “.txt”, etc. As summarized in Figure 8, TSLib currently supports more than 30 datasets with different sampled frequencies across four mainstream time series analysis tasks, all derived from real-world scenarios, such as energy, transportation, economics, weather, and medicine etc. Moreover, TSLib excels in scalability, allowing for the effortless integration of new data sources of different types. Besides, to support subsequent training and evaluation, TSLib includes a multitude of data processing steps, including time window splitting, data batch generation, etc, where the raw data is partitioned into separate windows for training, validation, and testing purposes, enabling streamlined model training and equitable comparisons. These steps serve as indispensable prerequisites for attaining precise and dependable results across a range of diverse time series analysis tasks.
@@
Moreover, TSLib provides additional support for numerous crucial and effective data processing strategies [36, 37, 38, 43] to enhance model performance and training efficiency. We encapsulate these various design strategies as basic data processing layers, encompassing techniques such as data normalization, time-frequency decomposition, Fourier analysis, and more. When utilizing TSLib, users have the flexibility to select these strategies to improve the training effect based on their specific requirements and objectives.
Flexible Model Construction As stated in the previous sections, we abstract the construction of deep time series models as the combination of basic modules (Section 2) and model architectures (Section 4). Specifically, we separate the most commonly used modules, such as stationarization and decomposition, from previous elaborately designed models. In this way, it is very convenient to reproduce or add a deep model with pre-defined basic modules. Further, although different models are originally proposed for different tasks, such as Autoformer [22] is for forecasting and TimesNet [3] is for more generic tasks, following the “model-centric” principle in this survey, we unify diverse analysis tasks as a representation learning task with domain-specific heads, enabling a task-general evaluation for different backbones.
Comprehensive Performance Evaluation Building upon the above two essential features, TSLib can naturally support the evaluation of diverse models on different tasks based on extensive datasets, building a comprehensive benchmark. Specifically, in TSLib, we provide evaluation for four mainstream time series analysis tasks: classification, imputation, forecasting, and anomaly detection. Each task involves specific evaluation metrics, enabling a comprehensive assessment of the performance of models. Noticing that, we further distinguish the forecasting task as long- and short-term tasks following the previous convention, where the separation depends on the forecasting horizon.
5.2 Evaluation Protocols and Metrics
To conduct a fair and comprehensive model comparison, TSLib provides standardized evaluation protocols for four mainstream time series analysis tasks following [3]. Specifically, different models are trained and tested based on the same training strategies and test sets to avoid potential unaligned configurations and ensure a rigorous comparison among different models.